OpenStudy (anonymous):

simplify th following expression, and rewrite it in an equivalent form with positive exponents. -14x^4y^6/7x^3y^5. can someone please explain this one to me?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

is there anyone that can help me please

OpenStudy (anonymous):

I feel like a step-child

OpenStudy (anonymous):

well shucks !!!!!

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

-14xy/7 ?

OpenStudy (anonymous):

that is how you work the problem -14x^4y^6/7x^3y^5

OpenStudy (anonymous):

can you work this out to show me please

OpenStudy (anonymous):

i dint get you.. i thought that was the answer

OpenStudy (anonymous):

its x power 4 in the numerator and x power 3 in denominator .. this makes it x power 1 in the numerator ..similarly for y

OpenStudy (anonymous):

no that is the question. to simplify the following expression, and rewrite it in an equivalent form with positive exponents

OpenStudy (anonymous):

yes the question simplifies to -14xy/7

OpenStudy (anonymous):

can your write it all out to show me how this is worked out

OpenStudy (anonymous):

the choices are: 2xy or -2x^7y^11 or -2xy or 14xy

OpenStudy (anonymous):

-(14x^4 y^6)/(7x^3 y^5 )=-14xy/7=-2xy

OpenStudy (anonymous):

how did you get this work it out step by step please so i understand

OpenStudy (anonymous):

-(14x^4 y^6)/(7x^3 y^5 )=-(14x^((4-3) ) y^((6-5) ))/7=-14xy/7=-2xy

OpenStudy (anonymous):

Thank you very much!!!!!

OpenStudy (anonymous):

can you help me with one more problem

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yes i shall try

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Raise the quantity in parentheses to the indicated exponent,and simplify the resulting expression. (-5x^3y)^2

OpenStudy (anonymous):

(-5x^3y)^2=(-5)^2((x^3)^2)y^2=25x^6y^2

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

short answer. simplify the following expression, and rewrite it in an equivalent form with positiveecponents (-2x^-2)^-4

OpenStudy (anonymous):

i do not get these

OpenStudy (anonymous):

(-2x^-2)^-4=(-2)^-4(x^-2)^-4=(-2)-4x^(-2*-4)=x^8/(-2)4=x^8/16

OpenStudy (anonymous):

just concentrate on each terms and give each term their powers separately.. u can do it

OpenStudy (anonymous):

try doing it..once you understand how to give the powers it becomes simple..

OpenStudy (anonymous):

it should be -2^-4 and 2^4 i forgot the power symbol

OpenStudy (anonymous):

where do i put these in place of

OpenStudy (anonymous):

just rewrite the whole problem and solution please

OpenStudy (anonymous):

(-2x^-2)^-4=(-2)^-4(x^-2)^-4=(-2)^-4x^(-2*-4)=x^8/(-2)^4=x^8/16

OpenStudy (anonymous):

thank you again

OpenStudy (anonymous):

just concentrate and you can do all these sums

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)